The mean speed of the molecules of an ideal gas is 2.0×10^3 ms^-1 . The radius of a gas molecule is 1.5×10^-10 m . Calculate the (i) collision frequency, and (ii) mean free path. It is given that 4×10^24 m^-3 .


1
Expert's answer
2021-03-25T14:08:33-0400

Answer

Mean free path

"\lambda=\frac{1}{\sqrt{2}\pi d^2 n}"

n="4\u00d710^{24 }m^{-3}"

d="3\u00d710^{-10}m"

v="2.0\u00d710^3 ms^{-1}"

"\lambda=\frac{1}{1.414\times3.14 ( 3×10^{-10}) ^2\times 4×10^{24 }m^{-3} }"

"\lambda=6256\times10^{-10}m"


Now collision frequency

"\nu=\frac{v}{\lambda}\\=\frac{ 2.0×10^3 }{ 6256\times10^{-10} }\\=31.98\times10^8Hz"